### v2

# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

Abstract.
The Smart Grid (SG) is a critical energy infrastructure that collects real-time electricity usage data to forecast future energy demands using information and communication technologies (ICT). Due to growing concerns about data security and privacy in SGs, federated learning (FL) has emerged as a promising training framework. FL offers a balance between privacy, efficiency, and accuracy in SGs by enabling collaborative model training without sharing private data from IoT devices. In this survey, we thoroughly review recent advancements in designing FL-based SG systems across three stages: generation, transmission and distribution, and consumption. Additionally, we explore potential vulnerabilities that may arise when implementing FL in these stages. Finally, we discuss the gap between state-of-the-art FL research and its practical applications in SGs and propose future research directions. These focus on potential attack and defense strategies for FL-based SG systems and the need to build a robust FL-based SG infrastructure. Unlike traditional surveys that address security issues in centralized machine learning methods for SG systems, this survey specifically examines the applications and security concerns in FL-based SG systems for the first time. Our aim is to inspire further research into applications and improvements in the robustness of FL-based SG systems.
Smart Grid, Federated Learning, Adversarial Attack, Privacy
copyright: acmlicensedisbn: ccs: Security and privacy Distributed systems security
1. Introduction
The Smart Grid (SG), which integrates the physical infrastructure of the electrical grid with modern information and communication technologies (ICTs) (vasisht2022smart), is a prime example of a cyber-physical system (CPS). Unlike conventional power grids, SGs can detect, respond to, and proactively manage changes in electricity demand and supply, while also addressing various operational challenges (obeagu2022overview).
The SG architecture incorporates several key components that work together to enhance the overall functionality of the electrical system (mittal2024smart), including advanced metering infrastructure (AMI) (molokomme2019survey), communication networks (hu2023multi), and sophisticated information technology solutions (reza2014overview), among others. By leveraging real-time data and advanced analytics, SGs aim to improve the efficiency of power generation, transmission and distribution, and consumption (chopda2021overview). Although SGs enable a more flexible and resilient power system that can better accommodate the growing demand for electricity, the increasing reliance on ICTs has made SG systems more vulnerable to security and privacy attacks. Federated Learning (FL) has emerged as a promising approach to addressing privacy and security concerns in SG systems. This innovative distributed machine learning paradigm enables collaborative model training across multiple participants without the need to share raw data, making it particularly suitable for SG applications, where generators are distributed and data privacy for each generator is crucial. We categorize the application of FL in SGs into three aspects: generation, transmission and distribution, and consumption. From the perspective of generation (tuballa2016review) in SGs, FL is primarily being used to enhance renewable energy integration and forecasting (camacho2011control). FL enables distributed SG generators to collaboratively train models for predicting resources such as solar and wind power output without sharing sensitive operational data. In the transmission (bayoumi2015power) and distribution aspects (hamidi2010smart) of SGs, FL is proving valuable for line fault detection and classification, optimal power flow (usman2022novel), voltage control (xie2021review), and more. By utilizing a FL-based approach, grid operators can quickly identify and locate faults in transmission lines, while collaboratively training models for power distribution without compromising the privacy of individual generators. In the consumption process of SGs (kaur2020energy), FL is extensively used for load forecasting (husnoo2023secure) and energy theft detection (li2024fuse). Smart meters can collaboratively train models to predict household energy consumption patterns without sharing individual usage data.
Table 1. Compared with State-of-the-art Survey Papers.
Ref Year Stage Applications DL-based Privacy Security FL-based
Grataloup et al. (grataloup2024review) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zheng et al. (zheng2024advancing) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zhang et al. (zhang2024vulnerability) 2024 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Pandya et al. (pandya2023federated) 2023 $\surd$ $\surd$ $\surd$ $\surd$
Cheng et al. (cheng2022review) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Mirzaee et al. (mirzaee2022smart) 2022 $\surd$ $\surd$ $\surd$
Ding et al. (ding2022cyber) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Cui et al. (cui2020detecting) 2020 $\surd$ $\surd$ $\surd$ $\surd$
Ours - $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
1.1. Related Surveys
We summarize the state-of-the-art (SOTA) survey papers focus on SGs, with particular attention to seven aspects: whether the paper discusses the stages of SGs, compares SG applications, includes Deep Learning (DL)-based methods, discusses privacy and security issues, and whether it focuses on FL-based SG systems (see Table 1).
Grataloup et al. (grataloup2024review) focus on the integration of FL to address challenges in renewable energy applications, such as energy management, system reliability, energy consumption prediction, and load forecasting. However, this survey only highlights FL use cases in SGs as privacy-preserving collaborative learning approaches and overlooks the potential vulnerabilities of SGs when using deep learning and federated learning technologies. Zheng et al. (zheng2024advancing) provide a detailed discussion of components in FL systems, including model aggregation, data partitioning, and local model training. However, their survey paper lacks a discussion on model-related attacks within FL-based SG systems. Zhang et al. (zhang2024vulnerability) offer a comprehensive survey of applications for each stage of SG systems and discuss potential privacy and security violations. However, their focus is primarily on centralized machine learning algorithms. Our work bridges the gaps between SG, FL, and privacy and security vulnerabilities. Pandya et al. (pandya2023federated) demonstrate a broader scope, covering smart cities with a FL framework to preserve privacy. This approach enables effective utilization of large-scale data in areas such as electric supply chain monitoring, air quality forecasting, water consumption prediction, traffic management, and pandemic diagnosis. However, this work lacks details on SG applications and security issues within FL frameworks. Cheng et al. (cheng2022review) conduct a survey of the FL paradigm, comparing the types, network structures, and global model aggregation methods in SG. However, the brief survey lacks a discussion on robustness. Mirzaee et al. (mirzaee2022smart) and Ding et al. (ding2022cyber) discuss conventional security threats to SGs as well as emerging issues related to the application of machine learning techniques. They cover various types of cyberattacks, privacy concerns related to smart meter data, and potential countermeasures to address these challenges. However, they only cover a limited number of applications in SGs and primarily focus on centralized cases. Cui et al. (cui2020detecting) mainly focus on false data attacks and privacy issues in SGs. Compared to previous literature, our paper provides a comprehensive discussion on FL-based SGs and their potential vulnerabilities. We hope this can complement existing research and inspire further studies in this direction.
1.2. Key Contributions
  1. (1)
    We provide a systematic study of existing research on FL-based SG applications, privacy issues, and security concerns. By defining and categorizing these into three stages (i.e., generation, transmission and distribution, and consumption) of SG applications, and two categories (i.e., privacy and security) of vulnerabilities, our holistic survey offers a hierarchical understanding of FL’s usage in SGs to preserve data privacy.
  2. (2)
    We compare our paper with other SOTA survey papers across seven aspects: whether the paper discusses the stages of SG applications, compares SG applications, includes DL-based methods, discusses privacy and security issues, and focuses on FL-based SG systems. This highlights the focus of our survey, which centers on FL-based SG applications and their potential vulnerabilities.
  3. (3)
    We discuss the use cases of the FL framework in three stages of SG systems. For the generation stage, we cover tasks such as energy forecasting, energy optimization, and energy management. In the transmission and distribution stage, FL has been applied to tasks such as energy distribution, attack detection, anomaly detection, and security assessment. In the consumption stage, FL can enhance tasks like optimal power usage, load forecasting, and energy theft detection. Incorporating FL at these stages has improved data privacy and distribution efficiency in SGs. After introducing FL applications to these stages, we discuss the potential vulnerabilities for FL-based SGs in both privacy and security contexts.
  4. (4)
    Our survey extensively discusses SOTA attack and defense methods in FL and identifies a significant gap between FL research and its applications in FL-based SG systems. We further suggest directions for future exploration in this domain.
1.3. Structure of the Paper
Figure 1. Structure of the Paper.
The structure of our paper is illustrated in Figure 1. In the following sections, we begin by exploring the foundational terminologies of SGs and the FL framework in Section 2. Next, Section 3 provides a comprehensive review of FL-based SG applications, detailing their implementation across various stages. In Section 4, we thoroughly examine the vulnerabilities in SG applications, covering both centralized machine learning and FL-based SG applications. Section 5 presents an overview of commonly employed attack and defense methods within general FL frameworks, identifies existing gaps, and proposes future research directions in the field. Finally, Section 6 concludes the paper.
2. Terminologies
2.1. Smart Grid
The electric power grid has evolved by leaps and bounds since its inception (masera2018smart). A major development in this domain was the integration of ICT. The integration of ICT with the traditional power grid facilitated distributed power generation via better resource management and automated control (simoes2012comparison). ICT integration also marked the evolution of the conventional grid into the SG and solved several limitations including unidirectional power flow and centralized bulk generation of electricity. Additionally, it reduced the power and energy domain’s dependence on fossil fuels like coal and natural gas, leading the way for sustainability and decarbonization via the development of microgrids (jahromi2023cyber; yadav2020microgrid). The SG power flow framework consists of the following stages:
  1. (1)
    Generation: This stage refers to the production of AC/DC power. In the conventional power grid, power generation was centralized in nature. However, with ICT integration for real-time monitoring, power generation can also be achieved in a regulated manner at residential buildings (for example using solar photovoltaic panels) and used for local consumption/fed back to the grid (kahrobaee2012multiagent).
  2. (2)
    Transmission/Distribution: This stage refers to the transfer of the power produced at the generation stage to the appropriate end nodes for utilization. This involves the use of transformers, transmission lines, substation equipment and other line devices like fault detector relays, breakers to ensure a safe and reliable transfer at appropriate bus voltage levels. The fault detector relays and breakers are often digital in nature and controlled by the central monitoring station via ICT devices (gaspar2023smart).
  3. (3)
    Consumption: This is final stage of the SG power flow process and involves the utilization of the power as received from the transmission and distribution networks for powering electrical devices. The consumption level can involve entities like government agencies, private industries, private individuals, etc., each of which can have a different consumption footprint. In the modern SG, integration of ICT devices means that some consumer-level devices like electric vehicles can also be used to transfer surplus/unused power back to the grid (coordinated via V2G communications (madawala2011bidirectional)).
While ICT integration in the power grid enabled easier and more robust resource management from a single centralized Supervisory Control and Data Acquisition (SCADA)-based control system (gaushell1993scada), the increasing number of IoT devices and local generation facilities like solar panels means that it can be extremely computationally resource intensive to monitor them. Inadequate monitoring facilities can mean that intrusive events like cyberattacks and line faults may not be detected and mitigated on time.
Slowed mitigation of cyberattacks, faults and other control emergencies can have a devastating impact on the stability of the SG (jahromi2023cyber). Moreover, the SG being a cyber-physical system implies that cyber layer vulnerabilities can be exploited to cause severe physical ramifications. For example, cyberattacks on the SG can lead to uncontrolled voltage increase at one or more buses leading to equipment damage and consumer-level device burnout (rath2022behind). To alleviate these issues, several researchers have proposed, implemented and evaluated the use of Artificial Intelligence (AI)-based tools and techniques for applications like real-time control (bose2017artificial), cybersecurity (ruan2023super), load forecasting (zhou2017holographic), and fault detection (sun2023high). These AI tools are typically deployed at the centralized controller level for monitoring, removing the need for human system operators.
As the AI modules employed for applications like cybersecurity/control monitoring and load forecasting in the SG environment, are often placed at a central server, they require the exchange of data between local Distributed Energy Resources (DERs) and the central node via communication networks. However, this presents privacy risks (sun2021solar) as the data being transferred can be utilized to reveal energy consumption patterns of private individuals (liu2018practical; zhao2020smart; wang2020spds). This is a major risk as the consumption patterns can be cross-analyzed to determine when the individuals are at home/outside. Further, their home appliances can also be revealed motivating malicious entities to schedule burglary attempts (liu2018practical).
Another risk that the communication-based training data exchange between local DERs and the centralized server housing AI modules is the non-willingness within third-party DER owners to share proprietary data. This non-willingness typically stems from the concern that the proprietary data may be accessed by malicious entities who can then know the system details of the DER. This can be used to manipulate the system via stealth cyberattack vectors (leng2021stability). Non-willingness can also arise from a lack of trust between the DER owners and the utility company that manages the SG and the control center.
Both the above-mentioned factors can lead to data constraints that may severely hinder the AI training process leading to flawed and untrustworthy decision-making, raising concerns on the reliability of AI. To solve the problem, An AI training strategy called Federated Learning can be used. In this strategy proprietary training data does not need to be transferred over a communication channel leading to alleviation of privacy concerns (mcmahan2017communication). More details about this training strategy are available in the following subsection.
2.2. Federated Learning
Federated Learning (FL) aims to collaboratively train models across multiple decentralized data owners holding potentially sensitive local data in a privacy-preserving manner. The essence of FL lies in its ability to learn a shared model by aggregating locally computed updates, rather than directly accessing or sharing the raw data. This machine learning (ML) paradigm not only enhances privacy and security but also enables the utilization of distributedly owned data for model training, making it particularly suitable for mission-critical applications, such as healthcare and finance, where data privacy is of paramount concern.
The typical FL process involves several key steps. Initially, a global model is distributed to all participating FL clients from the FL server. Each client then trains the ML model on its local data to derive an updated local model. These model updates are subsequently sent back to the FL server, where they are aggregated to update the global model. This cycle is repeated until convergence or specific performance criteria are met. A widely used framework in FL is Federated Averaging (FedAvg), renowned for its ability to aggregate model updates with minimal communication overhead. This is particularly crucial in FL settings, where a potentially large number of participants may have limited communication bandwidth.
To achieve meaningful real-world impact, FL must address several intricate challenges, including non-independently and identically distributed (non-IID) data across clients, system heterogeneity, and scalability. Non-IID data can introduce biases into models by favoring the data distribution of specific participants. System heterogeneity, characterized by variations in computation and communication capabilities among client participants, can result in uneven contributions to the model training process. Additionally, scalability concerns arise as the number of participants increases, necessitating the development of efficient algorithms and infrastructure to manage the aggregation of updates and distribution of the global model. The development of FL continues to evolve, spurred by advancements in both theoretical research and practical applications across various industries. FL is anticipated to play a pivotal role in building AI solutions that prioritize user privacy and data sovereignty.
The development of FL continues to evolve, spurred by advancements in both theoretical research and practical applications across various industries. FL is anticipated to play a pivotal role in building AI solutions that prioritize user privacy and data sovereignty.
Consider a typical FL setup involving $m$ IoT devices (e.g., smart grid sensors) and a central smart grid server, where the devices collaborate to train a global model $\mathbf{w}\in\mathbb{R}^{p}$ with dimension $p$, coordinated by the central server. Assume each IoT device $j\in[m]$ holds a local private dataset $\mathcal{D}_{j}$. The goal of FL is to solve the following optimization problem:
(1) $\min_{\mathbf{w}\in\mathbb{R}^{p}}F(\mathbf{w}):=\frac{1}{m}\sum_{j\in[m]}F_{j}(\mathbf{w}),$
where $F_{j}(\mathbf{w}):=\mathbb{E}_{z\in\mathcal{D}_{j}}[\ell(\mathbf{w};z)]$ represents the local loss for device $j$, and $\ell(\mathbf{w};z)$ is the loss function depending on the model parameter $\mathbf{w}$ and a data point $z$ sampled from $\mathcal{D}_{j}$. Note that the datasets $\mathcal{D}_{j}$ and $\mathcal{D}_{k}$ may have different distributions for $j\neq k$.
To solve the above optimization problem, the system runs for $T$ rounds of FL training. Initially, the server stores the global model $\mathbf{w}^{0}$. In round $t$, the server randomly selects a subset $\mathcal{Q}^{t}$ of $q$ devices and sends them the latest global model $\mathbf{w}^{t}$. These selected devices update the global model $\mathbf{w}^{t}$ using their local data. For instance, in the widely-used Federated Averaging (FedAvg) (fl_original) algorithm, each selected device performs $\kappa$ iterations of stochastic gradient descent (SGD) to update the local model as follows:
(2) $\mathbf{w}_{j}^{t,r+1}=\mathbf{w}_{j}^{t,r}-\alpha\cdot\mathbf{g}_{j}^{t,r},\quad\forall r=0,\ldots,\kappa-1,$
where $\mathbf{w}_{j}^{t,r}$ represents the local model of device $j$ during its training, initialized as $\mathbf{w}_{j}^{t,0}=\mathbf{w}^{t}$. Here, $\alpha$ is the local learning rate, and $\mathbf{g}_{j}^{t,r}:=\frac{1}{B}\sum_{z\in\xi_{j}^{t,r}}\nabla\ell(\mathbf{w}_{j}^{t,r},z)$ is the stochastic gradient computed over a mini-batch $\xi_{j}^{t,r}$ of size $B$ sampled from $\mathcal{D}_{j}$. After local training, each device computes the model update $\Delta_{j}^{t}:=\mathbf{w}^{t}-\mathbf{w}_{j}^{t,\kappa}$ and sends it to the server. In round $t$, the server aggregates the updates and applies them to the global model as follows:
(3) $\mathbf{w}^{t+1}=\mathbf{w}^{t}-\frac{1}{q}\sum_{j\in\mathcal{Q}^{t}}\Delta_{j}^{t},$
where $q=|\mathcal{Q}^{t}|$ is the number of selected IoT devices in round $t$.
3. FL-based SG Applications
Figure 2. Stages of SG Systems. Electricity is produced by utilities and independent power producers at various types of power plants. The electric transmission system serves as a crucial connection between power generation and consumption. High-voltage transmission lines transport electricity across long distances from power plants to populated areas. At substations, the voltage is stepped down to lower levels, enabling distribution companies to deliver the power to local communities. Finally, the electricity is used by consumers for everyday purposes at the point of consumption.
The process of delivering electricity to consumers involves several key stages (as shown in Figure 2). First, electricity is generated at power plants by utilities and independent power producers using various energy sources such as fossil fuels, wind, and solar. Next, the transmission system carries the electricity over long distances through high-voltage lines from the generation sites to substations. At substations, the voltage is reduced to lower levels for safe distribution. The distribution network then delivers this electricity to homes, businesses, and other consumers. Finally, the electricity is used by consumers for daily activities, powering appliances, devices, and industrial equipment.
FL has emerged as a powerful framework for various SG applications, offering privacy-preserving and distributed solutions for tasks such as energy forecasting, energy optimization, energy management, energy distribution, attack detection, anomaly detection, security assessment, optimal power usage, load forecasting, and energy theft detection, among others. In the following subsections, we discuss how each of these works utilizes FL in SG applications.
3.1. Generation Stage
FL enables collaborative model training across different energy producers and consumers, enhancing the accuracy of energy generation forecasts and enabling dynamic resource allocation without the need to centralize sensitive operational data (lee2023towards; gupta2023fedgrid). In the context of energy management (abdellatif2024sdcl), FL supports the development of intelligent systems that can predict and optimize power generation patterns while keeping sensitive user data local on smart devices. This approach improves grid efficiency and facilitates more effective demand response programs.
Figure 3. The importance of energy optimization in complex resource and consumption scenarios.
Energy Forecasting In the work by (lee2023towards), an advanced algorithm for FL is proposed to tackle the challenges of non-IID data in distributed solar energy grids. This framework is specifically designed to handle the heterogeneous nature of the data across different solar energy producers, enhancing the model’s ability to converge efficiently in decentralized settings. By addressing the variations in data distribution, the framework improves the accuracy and reliability of predictions, leading to better performance in solar energy generation and more effective resource allocation.
Energy Optimization The FedGrid framework introduced by (gupta2023fedgrid) offers a secure FL solution for optimizing energy management within smart grids. This framework emphasizes the importance of protecting privacy and enhancing data security during the optimization process. By enabling efficient collaboration between various grid components without compromising data confidentiality, FedGrid helps optimize energy distribution and usage. The approach also addresses the key challenges associated with data sharing and collaboration in smart grids, ultimately improving the overall efficiency and reliability of the system. As shown in Figure 3, energy optimization is crucial for coordinating multiple energy resources and consumption, especially when using FL techniques to protect data privacy.
Energy Management In (abdellatif2024sdcl), a Secure, Distributed, and Collaborative Learning (SDCL) framework is presented to enhance the operations of smart grids. This framework focuses on maintaining data privacy and security while facilitating collaborative learning among grid components. By incorporating advanced AI and blockchain technologies, SDCL aims to optimize grid performance, improve system efficiency, and ensure reliable energy management. Additionally, (lin2024privacy) introduces a privacy-preserving energy management framework that utilizes vertical FL for smart data cleaning in smart homes. This approach enhances user privacy while enabling effective analysis and management of smart meter data, ensuring that energy usage is optimized without compromising sensitive information.
3.2. Transmission and Distribution Stages
For energy distribution (mourtzis2023integration), FL enables optimal energy distribution solutions and flexibility predictions, significantly improving the efficiency of power distribution in SGs. In the area of attack (tran2023efficient; zhao2024federated) and anomaly detection (kesici2024detection; jithish2023distributed), FL has demonstrated its effectiveness in identifying cyber attacks and anomalies within SG systems. This approach allows multiple stakeholders to collaboratively train robust detection models while maintaining data privacy, as raw consumption data remains local. In the realm of assessment (li2024privacy; ren2023qfdsa; ren2023efeddsa), FL has been applied to dynamic security assessment, enabling participants to collaboratively build models while preserving data privacy.
Energy Distribution Mourtzis et al. (mourtzis2023integration) develop an integrated FL framework aims at enhancing the efficiency and security of energy distribution in smart grids. This framework addresses privacy concerns by enabling collaborative learning without sharing raw data, which is particularly relevant in smart grid data analysis. Specifically, they propose an incentive mechanism for FL, edge devices are compensated by service providers based on their contribution to a given task. The payoff function for an edge device $i$ participating in task $\Upsilon_{j,k}$ under service provider $j$ is:
(4) $u_{i,j,k}(Q_{i,j,k},S_{i,j,k})=\varpi\kappa_{p}p_{i,j,k}\Psi_{i,j,k}-(1-\varpi)C_{i,j,k},$
where $\varpi$ is a weight parameter that balances revenue and cost, $\kappa_{p}$ is an adjustment parameter, $p_{i,j,k}$ is the payment to edge device $i$, and $\Psi_{i,j,k}$ is the highest Age of Local Model achieved by $i$ in task $\Upsilon_{j,k}$. The total cost $C_{i,j,k}$ for edge device $i$ includes local data cost, computational cost, and communication cost. The payoff function for service provider $j$ in task $k$ is given by:
(5) $\pi_{j,k}=\mu\nu_{a}\Lambda_{j,k}-(1-\mu)\left(\sum_{i\in\mathcal{I}_{j,k}}p_{i,j,k}\Psi_{i,j,k}+\lambda_{c}\Xi^{\text{comp}}_{j,k}\right),$
where $\Lambda_{j,k}$ represents the accuracy of the global model for task $\Upsilon_{j,k}$, and $\Xi^{\text{comp}}_{j,k}$ is the cloud aggregation cost. $\mu$ is a weight parameter that balances model accuracy and total payment.
Attack Detection Tran et al. (tran2023efficient) introduce a privacy-preserving FL approach specifically designed to detect false data injection attacks (FDIAs) in smart grid systems. This method focuses on enhancing data privacy and security in cross-silo FL scenarios within the SG environment. By improving the detection efficiency and effectiveness of FDIAs, their approach ensures that data confidentiality is maintained across various entities involved in the grid, making it a robust solution for securing SG infrastructure. Zhao et al. (zhao2024federated) propose a FL algorithm that integrates client sampling and gradient projection techniques to enhance SG applications. This approach effectively addresses privacy concerns by enabling model training using fragmented data from devices spread across the grid. Furthermore, the inclusion of client sampling and gradient projection techniques leads to improved FL performance and optimization, providing a more reliable framework for handling SG data without compromising security.
Anomaly Detection Kesici et al. (kesici2024detection) propose a data-driven time series anomaly detection method that leverages time-aware shapelets, a shapelet evolution graph, and segment embeddings learned through the DeepWalk algorithm. Their method demonstrates a significant improvement in detecting FDIA, with a 20-40% increase in accuracy compared to other unsupervised learning techniques. Additionally, when applied with BO-XGBoost, the model achieves a 5% higher accuracy than unmodified XGBoost, adding further robustness to the detection process. Moreover, this approach imparts physical significance to the dynamic evolution of time series models, providing valuable insights into the underlying data. Jithish et al. (jithish2023distributed) present a FL-based approach for distributed anomaly detection in SGs. Their solution enhances privacy and security by allowing for collaborative detection of anomalies without the need to centralize sensitive data. This distributed approach not only safeguards data privacy but also improves the overall stability and reliability of SG systems. By enabling effective anomaly detection across the grid, their method ensures that potential issues are identified early, contributing to more resilient SG operations.
Security Assessment Li et al. (li2024privacy) introduce a quantum-secured FL system, known as QFDSA, for dynamic security assessment in SGs. This system integrates quantum-resistant cryptography with FL to ensure both data privacy and security, particularly in the context of SG dynamic security assessments (DSA). Their approach addresses the limitations of traditional DSA methods, which may be vulnerable to privacy breaches and quantum computing attacks, thereby offering a more secure alternative for assessing grid security. Ren et al. (ren2023qfdsa) further develop QFDSA, a hybrid system that combines quantum security and FL for dynamic security assessment in SGs. By enhancing privacy and security measures, their approach tackles key challenges related to data protection and collaborative analysis in SG environments. Specifically, in Quantum Key Distribution (QKD) for FL, secure keys are exchanged between participants, typically denoted as Alice and Bob, using quantum states to ensure security. In a Measurement-Device-Independent QKD (MDI-QKD) protocol, both Alice and Bob randomly choose a measurement basis from either $K$ or $Z$ and send weak coherent-state pulses (WCPs) with intensities sampled from predefined sets. These pulses are transmitted to a trusted or untrusted relay (Charlie) which performs a Bell state measurement (BSM). The key rate $R$ for secure communication is calculated by:
(6) $R=P_{s}^{A}P_{s}^{B}\left[s_{A}e^{-s_{A}}\cdot s_{B}e^{-s_{B}}\cdot Y_{11}^{K,L}\left(1-h(e_{11}^{K,U})\right)-feQ^{Z}h(E^{Z})\right].$
Here, $P_{s}^{A}$ and $P_{s}^{B}$ resent the probabilities that Alice and Bob send WCPs with the intensity of the signal state. $Y_{11}^{K,L}$ and $e_{11}^{K,U}$ are the yield and error rates in the decoy basis $K$, respectively. $f$ is the error correction efficiency, and $Q^{Z}$ and $E^{Z}$ are the gain and error rates in the signal basis $Z$, ensuring quantum-secured keys for FL.
Ren et al. (ren2023efeddsa) propose EFedDSA, an efficient differential privacy (DP)-based horizontal FL approach for SG dynamic security assessment. This system enhances privacy protection in data-driven security assessments, addressing the challenge of data isolation and privacy concerns in traditional machine learning approaches. EFedDSA improves the efficiency and effectiveness of dynamic security assessment while maintaining data confidentiality across multiple participants in the SG infrastructure, offering a comprehensive solution for modern SG security needs.
3.3. Consumption Stage
In the area of optimal power usage (rajesh2024federated; abdulla2024smart), FL facilitates the extraction of typical electricity consumption patterns from smart meter data, enabling more efficient grid management. For load forecasting (badr2023privacy; liu2023fedforecast; fekri2023asynchronous), FL-based methods have been developed to enhance the accuracy of short-term residential load predictions by utilizing distributed data from multiple sources. In the case of electricity theft detection (li2024fuse; alshehri2024deep; zafar2023step), frameworks like FedDetect employ federated learning to effectively identify fraudulent activities while preserving consumer privacy.
Optimal Power Usage Rajesh et al. (rajesh2024federated) propose a FL framework designed to enhance privacy while optimizing power usage in SGs. The framework allows for collaborative learning across multiple participants without compromising individual user data. By securing power traces and enabling personalized recommendations for power usage optimization, the approach significantly improves privacy preservation in SG data analytics. This personalized aspect ensures that recommendations are tailored to the specific needs of users while maintaining data confidentiality. Similarly, Abdulla et al. (abdulla2024smart) present an adaptive FL approach for energy consumption forecasting using smart meter data. Their method improves forecasting accuracy by utilizing smart meter data without sharing raw information, ensuring user privacy. The adaptive nature of the approach allows it to adjust to individual consumption patterns, providing personalized forecasting models for each participant. This enhances both the accuracy of energy predictions and the privacy of users in the SG system.
Load Forecasting Badr et al. (badr2023privacy) introduce a privacy-preserving and communication-efficient energy prediction scheme based on FL for SGs. Their approach enables accurate load forecasting while maintaining data confidentiality, which is crucial for infrastructure planning and power dispatching. By enhancing communication efficiency in the FL process, this method helps reduce power outages and equipment failures, making it an important tool for SG operations. Liu et al. (liu2023fedforecast) propose FedForecast, a FL framework designed for short-term probabilistic individual load forecasting. This framework provides comprehensive information about future load uncertainties while addressing data isolation and privacy concerns inherent in traditional centralized forecasting methods. By maintaining data confidentiality across households or smart meters, FedForecast improves the accuracy and reliability of short-term load forecasts. Fekri et al. (fekri2023asynchronous) develop an asynchronous adaptive FL algorithm for distributed load forecasting with smart meter data. This approach enhances the scalability and efficiency of load forecasting in SGs by adapting asynchronously to the data, thus addressing issues of privacy, security, and communication overhead found in traditional centralized methods. By doing so, the algorithm ensures accurate load forecasting while preserving the privacy of distributed data.
Energy Theft Detection Li et al. (li2024fuse) introduce FUSE, a framework that combines FL with U-shape split learning for detecting electricity theft. FUSE identifies abnormal electricity consumption behaviors while preserving the privacy of both user data and the detection model. By integrating FL with split learning, FUSE enhances data protection and model security in smart grid systems, offering a robust solution for electricity theft detection. Alshehri et al. (alshehri2024deep) develop a deep anomaly detection framework that utilizes FL to detect electricity theft, including zero-day threats, in smart power grids. This framework preserves privacy while effectively identifying new and previously unseen attacks, thereby enhancing the security and stability of SGs. By addressing the challenge of detecting sophisticated electricity theft attempts, the framework plays a critical role in maintaining grid reliability. Zafar et al. (zafar2023step) present a FL-assisted hybrid deep learning model for electricity theft detection using smart meter data. This model enhances the security and reliability of SGs, particularly in the context of Industry 5.0. By leveraging FL to keep data decentralized, the model addresses privacy concerns while effectively detecting electricity theft. This hybrid approach offers a comprehensive solution to safeguard SGs from fraudulent activities.
4. Vulnerabilities in FL-based SG Systems
Figure 4. The classification of vulnerabilities in FL-based SG systems.
In this section, we explore the vulnerabilities present in SG systems. As illustrated in Table 2, our discussion is structured around three critical stages within SGs: generation, transmission and distribution, and consumption, as well as two vulnerability aspects, as depicted in Figure 4: privacy and security. This systematic approach allows us to comprehensively demonstrate the potential vulnerabilities associated with various applications within these stages. In each subsection, In each subsection, We begin by presenting the fundamental concepts underlying each application to establish a clear understanding of their functionalities and objectives. We then examine the specific aspects that make these applications potentially vulnerable, highlighting their inherent risks.
4.1. Vulnerabilities in Generation Stage
In the context of SG, generation refers to the production of electricity from various sources, including both traditional and renewable energy (tuballa2016review). It is a crucial component of the SG pipeline, serving as the starting point for the flow of electricity throughout the entire system. As generation systems become more integrated and digitalized, privacy and security becomes increasingly important (krause2021cybersecurity). Protecting generation infrastructure from attacks is essential for maintaining the reliability and integrity of the entire SG system.
4.1.1. Energy Forecasting
A key aspect of generation in SGs is the integration of energy forecasting, particularity for renewable energy sources. SGs are designed to accommodate and efficiently manage the intermittent nature of renewable energy generation, such as solar and wind power (wang2023investigation). This integration helps reduce reliance on fossil fuels and promotes a more sustainable energy future. Specifically, energy forecasting leverages advanced machine learning techniques, including deep learning models like Long Short-Term Memory (LSTM) networks, to analyze historical data and predict future energy generation patterns (malik2022integration).
Although accurate forecasting helps maintain grid stability by predicting fluctuations in renewable energy output, it also introduces new vulnerabilities, as attackers may target the prediction process. In the context of the FL framework, Shabbir et al. (shabbir2024resilience) focus on energy forecasting using FL in SG applications. The data involved typically consists of time-series data related to energy consumption and generation patterns. one primary attack method discussed is FDIA, which can take various forms, such as scaling attacks, random noise attacks, and mean-based attacks. These attacks aim to manipulate the learning process by injecting falsified data into the FL system. The impact of these attacks is often measured using metrics like Mean Absolute Percentage Error (MAPE). For instance, experiments have shown that scaling attacks can increase MAPE from 0.193% to 32.72%, while random noise attacks can elevate it from 0.183% to 129.75% as the proportion of attacked data rises from 10% to 100%. Badr et al. (badr2022privacy) propose a privacy-preserving FL approach for net-energy forecasting, addressesing the critical need for accurate energy predictions while protecting individual privacy in SG systems. To enhance privacy protection, researchers have introduced various strategies, including secure data aggregation (fereidooni2021safelearn), differential privacy, and hybrid mechanisms. These approaches aim to mitigate potential privacy leaks that may occur through attacks on shared model parameters. FedGrid (gupta2023fedgrid) is an innovative framework that leverages FL to optimize energy distribution in SG systems. It addresses two major challenges in the energy sector: accurately predicting renewable energy generation and forecasting electric load, while simultaneously maintaining data privacy and security. By utilizing FL, FedGrid enables collaborative model training across multiple parties without compromising the privacy of individual data, a crucial factor when handling sensitive information like user electricity consumption patterns and localized weather data. FedGrid’s comprehensive analysis helps SG operators develop efficient supply strategies, optimizing the distribution of energy from both renewable and non-renewable sources.
Table 2. Paper Reviews on Vulnerabilities in SG Systems.
Stage Ref Year Task Data Type Attack Attack Type
Generation Shabbir et al. (shabbir2024resilience) 2024 Energy Forecasting Time Series False Data Injection Security
Badr et al. (badr2022privacy) 2022 Net-Energy Forecasting Time Series Model Inversion, Membership Inference Privacy
FedGrid (gupta2023fedgrid) 2023 Net-Energy Forecasting Miscellaneous Data Inference Privacy
Moayyed et al. (moayyed2022cyber) 2022 Wind Speed Forecasting Miscellaneous Data Inference Privacy
T&D Lin et al. (lin2022incentive) 2022 State Estimation Measurements False Data Injection Security
Mohammadabadi et al. (mohammadabadi2023towards) 2023 Event Classification Phasor Measurement Unit Data Inference Privacy
Jithish et al. (jithish2023distributed) 2023 Anomaly Detection Miscellaneous Data Integrity Security
Liu et al. (liu2022federated) 2022 Voltage Control Miscellaneous Data Inference Privacy
Consumption Husnoo et al. (husnoo2023secure) 2023 Load Forecasting Historical Data Integrity Security
Wen et al. (wen2021feddetect) 2021 Energy Theft Detection Irish Smart Meter Inference, Model Poisoning Privacy, Security
Ashraf et al. (ashraf2022feddp) 2022 Energy Theft Detection Smart Meter Data Poisoning Security
Zafar et al. (zafar2023step) 2023 Energy Theft Detection Irish Smart Meter Data Poisoning Security
Li et al. (li2024fuse) 2024 Energy Theft Detection Smart Meter Data Poisoning, Model Inversion Privacy, Security
4.1.2. Wind Speed Forecasting
Wind speed forecasting is a crucial process in renewable energy management, involving the prediction of future wind speeds using advanced statistical and machine learning techniques (alhussein2019microgrid). This method utilizes historical wind speed data, along with other meteorological parameters, to generate accurate predictions for various time horizons, ranging from short-term (hours ahead) to long-term (days or weeks ahead) (malik2022multi). Accurate wind speed forecasting directly contributes to improved wind power prediction, enabling the seamless integration of wind energy into the SG. By anticipating wind power output, grid operators can better optimize the balance between wind energy and other power sources, resulting in more efficient utilization of renewable resources (ghofrani2014smart).
Wind speed forecasting often relies on advanced machine learning and deep learning algorithms (lin2024prediction; sun2021solar), which can be vulnerable to adversarial attacks designed to manipulate their outputs. Attackers could exploit weaknesses in these algorithms, introducing errors in wind speed predictions and leading to inaccurate power generation forecasts. Within the FL framework, Moayyed et al. (moayyed2022cyber) propose a wind power forecasting framework integrated into the FL pipeline. The system predicts future wind power generation using multiple resource, including time-series data on wind power generation and meteorological data. Additionally, the model incorporates Sentinel-1 and Sentinel-2 satellite images for wind speed and bathymetry analysis, enhancing the spatial information available for forecasting. While the FL approach inherently provides some defense against data breaches by decentralizing sensitive information, local model updates from attackers can potentially compromise the entire system (zhang2015byzantine; liu2020reflection).
4.2. Vulnerabilities in Transmission and Distribution Stages
SG technology revolutionizes traditional power grids by incorporating advanced ICTs into the existing energy infrastructure. This integration enables a more efficient, reliable, and flexible system for electricity transmission and distribution (hamidi2010smart). The transmission component of the SG (bayoumi2015power) focuses on the high-voltage transport of electricity over long distances. In SGs, transmission systems benefit from advanced technologies such as Flexible AC Transmission Systems (FACTS) and High-Voltage Direct Current (HVDC) power transmission, which enhance the controllability, stability, and power transfer capability of AC transmission systems. However, transmission systems are vulnerable to various types of attacks. FDIAs are particularly concerning, as they can manipulate sensor measurements and mislead grid operators.
SG distribution networks (konstantelos2016strategic) are responsible for delivering electricity to end-users at lower voltages. These networks are undergoing significant changes to accommodate the increasing integration of DERs and renewable energy sources. Smart distribution systems enable bi-directional power flow, allowing consumers to also become producers of electricity. However, distribution systems face threats such as energy theft through meter tampering, privacy breaches of consumer data, and coordinated attacks on demand response systems (ebtia2024spatial).
4.2.1. State Estimation
AC State Estimation is a crucial process in SG systems, designed to determine the most likely steady-state operation of the power system based on field measurements. It serves as a fundamental data processing tool responsible for supporting and increasing system visibility while filtering errors that may appear in real-time measurements, system topology, and parameters (acurio2023state).
However, despite its significance, AC State Estimation is vulnerable to cyber threats, particularly FDIAs. These attacks can manipulate measurement data, introduce errors in state estimation, which can impact bus power flow calculations and the economic dispatch of the power system (huang2022false). Within the FL framework, Lin et al. (lin2022incentive) propose an edge-based FL framework for detecting FDIAs in SG state estimation. This approach leverages local system operating data to predict and assess system stability in a distributed manner. By utilizing edge devices for local model training, the framework reduces the computational burden on central servers and enhances privacy protection. A key innovation in this work is the incorporation of an incentive mechanism designed to encourage high-quality data sharing and model training. This mechanism addresses the challenge of motivating participants to actively contribute to the FL process, which is essential for improving the overall detection accuracy of the system.
4.2.2. Events Classification and Anomaly Detection
Grid events classification involves identifying and categorizing various disturbances or anomalies that occur within power systems. This process is crucial for maintaining the reliability, stability, and security of electrical grids. By accurately detecting and classifying these events, power system operators can quickly respond to potential issues and mitigate their impacts on the grid operations.
Deep learning-based models (niazazari2020attack), which have gained popularity for grid event classification, are particularly vulnerable to adversarial attacks. While highly effective under normal conditions, these models can be deceived by carefully crafted perturbations in the input data. The vulnerability arises from the complex nature of these models and their sensitivity to small changes in input features. Within the FL framework, Mohammadabadi et al. (mohammadabadi2023towards) propose a federated XGBoost model to address data privacy concerns and enable collaborative learning. One key challenge in implementing FL for phasor measurement unit (PMU) data is the varying numbers of PMUs across different entities, resulting in diverse input structures for the event classification model. To overcome this, the proposed model constructs a common set of event features that can be generated by each entity, regardless of how many PMUs they possess. This study also considers two types of adversarial attacks on images using the Fast Gradient Sign Method (FGSM) and an iterative variant. In both approches, perturbations are applied to clean images to generate adversarial examples. For untargeted attacks, FGSM updates the adversarial image as follows:
(7) $x_{\text{adv}}=x+\epsilon\cdot\text{sign}(\nabla_{x}L(x,y)),$
where $\epsilon$ controls the perturbation size, $L(x,y)$ is the loss, and $\nabla_{x}$ represents the gradient. For targeted attacks, an iterative method is used:
(8) $x_{\text{adv}}=x+\Delta,$
where $\Delta_{i+1}=\Delta_{i}-\alpha\cdot\text{sign}(\nabla_{x}L(x,y))$ for each iteration $i$, with constraints $-\epsilon\leq\Delta_{i+1}\leq\epsilon$. These perturbations are visually subtle but malicious, creating adversarial images that fool machine learning models. To mitigate theses attacks, the paper employs data processing techniques and a LSTM network. To further strengthen privacy protection in the FL process, the paper incorporates an additive homomorphic encryption scheme. This encryption method secures the communication during the training process, eliminating the need for a trusted server and enhancing overall data protection.
Jithish et al (jithish2023distributed) propose a FL-based solution that enables the training of machine learning models across multiple distributed devices without compromising user privacy or relying on extensive centralized computing resources. This approach not only preserves data confidentiality but also enables real-time anomaly detection, which is crucial for identifying and responding to various grid issues such as unusual power consumption patterns, infrastructure faults, and potential attacks. In this system, a global model is downloaded from the server to smart meters for on-device training. After local training, only the updated local model parameters are sent to the server to improve the global model, thus ensuring user privacy. To further protect the model parameter updates from potential adversaries, the SSL/TLS protocol is employed. adding an extra layer of security to guarantee the integrity and confidentiality of the transmitted data.
4.2.3. Voltage Control
Voltage regulation is a critical process in electrical power systems, designed to maintain voltage levels within acceptable limits across the network. It involves controlling and adjusting the voltage to ensure stable and reliable power delivery to consumers (xie2021review). However, these networks are susceptible to data-related attacks, which can lead to voltage instability. Attackers may exploit vulnerabilities in communication channels to inject false data or disrupt the flow of information critical for effective voltage control.
In the FL-based SG system, Liu et al. (liu2022federated) combine FL with reinforcement learning techniques to develop a privacy-preserving and efficient control strategy. In this federated reinforcement learning framework, grid-connected interface inverters of renewable energy sources are modeled as deep neural network-based agents that interact with their local environment to learn optimal voltage-var control strategies. The process operates in a two-stage process: local training using reinforcement learning algorithms, followed by periodic global aggregation of model updates. This approach offers several advantages, including enhanced data privacy, reduced communication costs, improved scalability, and adaptability to changing network conditions.
4.3. Vulnerabilities in Consumption Stage
The consumption part of a SG represents the end-user side of the energy distribution system (kaur2020energy), encompassing homes, businesses, industries, and more. It integrates advanced technologies like AMI with smart meters for real-time usage monitoring and demand-side management to optimize energy consumption patterns. This system leverages IoT devices and real-time analytics to enhance energy efficiency and provide consumers with detailed insights into their usage. Additionally, sophisticated forecasting techniques and machine learning algorithms are employed to predict and manage energy loads effectively, ensuring a balance between supply and demand. However, smart meters and IoT devices, which are integral to these systems, can be vulnerable to attacks (liu2021vulnerability), potentially allowing malicious customers to manipulate energy consumption data or disrupt power supply.
4.3.1. Load Forecasting
Smart meter measurements, while essential for accurate demand forecasting, raise significant privacy and data breach concerns for consumers. The fine-grained load data collected by smart meters can potentially reveal sensitive information about residential users’ daily routines and behaviors, posing risks to property security (husnoo2023secure). In the FL framework, Husnoo et al. (husnoo2023secure) propose a secure FL approach designed for residential short-term load forecasting. This approach addresses the critical balance between accurate demand prediction and protecting consumer privacy in smart meter data. FL allows collaborative model training without exposing raw data, as models are trained locally, and only parameters are shared with a central server. To enhance security, the framework incorporates advanced techniques like gradient quantization using SignSGD (bernstein2018signsgd) and differential privacy (dwork2006differential). Additionally, it employs hierarchical and peer-to-peer strategies to improve both privacy and accuracy. The system tackles challenges such as non-IID data distribution and potential attacks through methods like deep federated adaptation and secure aggregation. Experimental results demonstrate that this approach achieves high forecasting accuracy while preserving user data privacy, with significant improvements in prediction accuracy compared to traditional methods.
4.3.2. Energy Theft Detection
Energy theft detection (salinas2015privacy) in SGs refers to the process of identifying and preventing unauthorized consumption of electricity, an illegal activity that causes significant economic losses to utility companies. SGs integrate digital technologies with traditional power systems, enabling advanced energy management and monitoring capabilities. These systems utilize AMI and smart meters to collect, measure, and analyze energy usage data from customers at regular intervals. However, the collection and analysis of consumer energy data create vulnerabilities, potentially allowing attackers to compromise smart meters.
Within the FL framework, Wen et al. (wen2021feddetect) introduce FedDetect, a novel privacy-preserving FL framework designed for energy theft detection in SGs. FedDetect leverages FL to train machine learning models on distributed devices, aggregating local model updates instead of raw data to protect user privacy. Ashraf et al. (ashraf2022feddp) propose FedDP, a privacy-protecting theft detection scheme for SGs using FL. This framework addresses the challenge of utilizing consumer energy data for theft detection while maintaining privacy protection. Zafar et al. (zafar2023step) develop a FL-assisted hybrid deep learning model for electricity theft detection in SGs. Their model combines FL techniques with deep learning to enhance security and reliability of electricity theft detection. Li et al. (li2024fuse) present FUSE, a novel electricity theft detection framework that integrates FL with U-shape split learning. FUSE aims to address the challenges of data privacy and security in SG systems while effectively detecting electricity theft.
5. Discussions
In the previous sections, we summarize the applications of FL in SGs and the potential vulnerabilities discussed in the existing literature. Most works still focus on attack methods commonly used in centralized machine learning frameworks, such as FDIAs, untargeted adversarial attacks, and inference attacks. However, we propose that, compared to the SOTA research, there remains a significant gap between FL-based attack methods and the attack strategies explored in FL-based SG applications. In the following subsection, we will introduce SOTA FL attacks (i.e., Byzantine attacks, backdoor attacks, and inference attacks) and SOTA FL defense mechanisms. Finally, we will discuss several potential future research directions.
5.1. SOTA FL Attack
5.1.1. Byzantine Attack in FL
Byzantine attacks (aka model poisoning attacks) in FL involve malicious clients sending incorrect or misleading model updates to the central server, aiming to disrupt the training process and degrade model performance. These malicious local model updates are typically intentionally modified by those malicious clients controlled by the attacker. We introduce several classical and commonly discussed Byzantine attacks as follows.
Random attack. To disturb the convergence of the global model, the malicious clients send randomized updates that follow a Gaussian distribution $N(\mu,\sigma^{2}\textbf{I}_{d})$ where $d$ is the dimension of the local model update. In particular, $\mu$ and $\sigma$ should be carefully chosen to make the malicious model updates have a similar distribution to benign ones.
Figure 5. The Explanation of the FL Framework and Its Potential Attack Points.
Noise attack. The malicious clients perturb their local model updates by adding Gaussian noise sampled from $N(\mu,\sigma^{2}\textbf{I}_{d})$. Here, $\mu$ and $\sigma$ should also be carefully chosen as a negligible noise (small $\mu$ and $\sigma$) cannot disrupt the convergence of the global model while a large noise (large $\mu$ and $\sigma$) makes malicious model updates significantly differ from benign ones.
Sign-flip attack. Instead of manipulating the magnitude of model updates, the malicious clients change the sign of their model updates to disturb the convergence of the global model. The Sign-flip attack is achieved by flipping the element-wise sign in the model updates.
Min-Max/Min-Sum attack (dnc). The creation of malicious model updates involves two steps. First, the attacker generates a malicious update by perturbing the average of all benign updates. In the second step, specifically for the Min-Max attack, the attacker optimizes this malicious update to ensure that its maximum Euclidean distance from any benign update is no greater than the maximum distance between any two benign updates. As for Min-Sum attack, the malicious update is optimized to ensure that the sum of its distances with each benign update is upper-bounded by the maximum total distance of a benign update among other benign updates. The Min-Max and Min-Sum attacks are powerful since malicious clients try to search for an optimal perturbation such that it has maximum impact.
AGR-tailored Trimmed-mean attack (dnc). In (dnc), the authors proposed several attack methods specific to some well-known defense methods. We take the AGR-tailored Trimmed-mean attack as a example. The AGR-tailored Trimmed-mean is designed to attack the defense method Trimmed-mean proposed in (trmean) by maximizing the Euclidean distance between the aggregated result of simple average and Trmean. This straightforward and intuitive method is effective, but the malicious model updates produced by the AGR-tailored Trimmed-mean attack differ significantly from benign ones, making them susceptible to easy detection by defense methods.
Lie attack (lie). In A-little-is-enough attack, the malicious clients apply slight perturbation to their local benign updates, making it hard to be detected by the defense methods. Specifically, the malicious clients calculate the element-wise mean $\mu_{j}$ and standard error $\sigma_{j}$ of all updates and generate the element of malicious updates by $(\beta_{i})_{j}=\mu_{j}-z\times\sigma_{j}$, where $\beta$ is the malicious model updates and $j\in[d]$. Here, $z$ is a scaling factor decided by the number of malicious clients and the total number of clients.
ByzMean attack (signguard). The ByzMean attack manipulates the mean of updates to be an arbitrarily malicious update. Specifically, it divides the malicious clients into two groups, with $m_{1}$ and $m_{2}$ clients respectively. The first group uses any existing attack methods (e.g., Lie attack) to generate their malicious updates. The second group creates their malicious updates to ensure that the overall average of all updates matches the average of the malicious updates from the first group.
5.1.2. Backdoor Attack in FL
Backdoor attacks (aka data poisoning attacks) in FL aim to maintain the global model’s performance on clean inputs while causing it to make incorrect predictions on inputs containing a specific predefined feature (also called trigger). Since backdoor attacks (lyu2024task; lyu2023attention) do not significantly affect the main task accuracy, the malicious local model updates are similar to benign ones, making anomaly detection more challenging in FL. We introduce several classical and commonly discussed backdoor attacks as follows.
Badnet attack (badnet). Badnet is a simple and effective backdoor attack where the attacker injects the same trigger into the training data of compromised clients. The ground truth label of these training data will also be changed to another backdoor class.
Distributed backdoor attack (dba). In the distributed backdoor attack (DBA), the adversary decomposes the backdoor trigger into several small triggers and applies them to different compromised clients rather than using the same trigger for all malicious clients (Badnet attack). This makes the attack harder to detect, as each client’s updates appear less suspicious.
Scaling attack (scaling). Based on Badnet attack, a scaling attack involves malicious clients scaling their model updates in a way that scales up the impact of their backdoor-triggered updates without affecting the clean data performance. The scaling factor ensures that the backdoor pattern has a stronger influence on the global model.
PGD attack (PGD). The Projected Gradient Descent (PGD) attack is adapted from adversarial machine learning. It iteratively perturbs the model updates of compromised clients within a certain bound to embed the backdoor. This method is effective at creating small but impactful changes to the model parameters.
Neurotoxin attack (Neurotoxin). Neurotoxin explores a persistent attack method designed for scenarios where attackers can only participate in a limited number of rounds. Specifically, in the Neurotoxin attack, the malicious clients project their model updates into a subspace that excludes the global coordinates. This approach ensures that the projected updates from malicious clients are primarily embedded in coordinates with minimal perturbation from benign updates, maintaining the backdoor effect even after the attack ceases.
5.1.3. Inference Attack in FL
Membership Inference Attack (MIA) (membership1; membership2) is a major privacy attack that seeks to determine whether a specific data record was part of the training dataset used to build a machine learning model. In the context of FL, MIA poses a significant threat to the privacy of participating clients and their data. Various entities, including malicious clients, servers, or external attackers, can execute MIA in FL. These adversaries exploit the model’s natural tendency to overfit on training data, leading to unintended information leakage. Approaches to conducting MIA in FL include attack models (membership3) and gradient-based MIAs, which exploit the gradients exchanged during the FL process. Passive MIA techniques also exist, where adversaries infer membership information by merely observing communication between clients and the server. Successful MIAs compromise the privacy of individual users, a particularly serious issue in sensitive domains such as healthcare or finance, where data confidentiality is paramount. Beyond identifying membership, MIA can also reveal attributes or properties of the training data, further threatening user privacy. Defending against MIA in FL presents unique challenges due to the distributed nature of the learning process, with traditional privacy-preserving techniques either being inapplicable or significantly impacting model utility. As a result, researchers are actively developing novel defense mechanisms to improve FL’s resilience to MIA while preserving model performance.
5.2. SOTA FL Defense
To defend against the attacks mentioned above, the server in FL usually adopts a defense method. Existing defense methods can be generally divided into three categories: detection-filtering methods (krum; bulyan; xu2022byzantine; LASA; signguard; dnc; fltrust), impact-reduction methods(trmean; bulyan; Foolsgold; lockdown; geomed; sparsefed; zhang2023byzantine), and differential privacy-based methods.
5.2.1. Detection-Filtering Methods
Detection-filtering methods aim to maximize robustness against attacks by identifying and removing malicious local model updates from the aggregation process. These methods typically rely on one or more statistical metrics (e.g., norms, cosine similarity, Euclidean distance) to assess local model updates. By comparing the extracted metrics with each update, they filter out updates that exhibit abnormal metric values, preventing them from influencing the global model. For example, Krum (krum) selects the most reliable local model that has the smallest sum of squared Euclidean distances to all other models as the output. Multi-Krum (krum) extends this by selecting multiple reliable models for averaging instead of just one. In (xu2022byzantine), the authors applied the KL-divergence to measure the distance between the empirical sign distributions of the local models in two consecutive rounds and used it to employ a weighted average of models for each client. The Divide-and-Conquer (DnC) method proposed in (dnc) calculates the product of a model with its principle component and removes the models with large projection values. FLTrust (fltrust) leverages a root dataset and trains a root global model based on it. After collecting local updates, FLTrust aggregates those updates with high Cosine similarity to root model updates. SignGuard (signguard) combines direction-based clustering and magnitude-based filtering to identify malicious model updates, where the direction and magnitude are obtained by the sign statistics and $L_{2}$-norm of the model update, respectively. Recently, a defense method called LASA that achieves SOTA results has been proposed in (LASA). It employs a layer-wise adaptive filter to adaptively select benign sparsified layers using both magnitude and direction metrics across all clients for aggregation.
5.2.2. Impact-Reduction Methods
Impact-reduction methods try to integrate all model updates but employ strategies to reduce the impact of malicious updates. For instance, Trimmed mean (Trmean) proposed in (trmean) discards a certain percentage of the highest and lowest values among the received models from clients for each dimension. After this trimming, the mean of the remaining values is computed by the server, which mitigates the impact of extreme values on the aggregated model. Mhamdi et al. (bulyan) proposed Bulyan that combines Trmean and Multi-Krum to extract benign models. In (geomed), the authors proposed to use the geometric median (GeoMed) of all local models as the aggregated model, which is effective in filtering out the outlier values. RLR (rlr) uses the sign information of model updates to adjust the global learning rate for each model coordinate without distinguishing between benign and malicious updates, resulting in poor accuracy. Foolsgold (Foolsgold) assumes that the malicious updates are consistent with each other. It assigns aggregation weights to model updates based on the maximum Cosine similarity between the last layers of pairwise model updates. A higher Cosine similarity value indicates a higher probability that the updates are malicious, leading to smaller aggregation weights being assigned. SparseFed proposed in (sparsefed) sparsifies the aggregated model update at the server side, integrating with model clipping and error feedback, to mitigate the impact of malicious local model updates. In addition, FedVRDP proposed in (zhang2023byzantine) applies sparsification on local model updates using a mask obtained from the global model in the previous round and combines it with random Gaussian noise perturbation and variance reduction techniques, to achieve robustness and differential privacy protection simultaneously. A recent approach called Lockdown (lockdown) assumes that those malicious parameters served to recognize backdoor triggers will be deemed unimportant for benign clients and applies sparsification to prune unimportant parameters on the local model.
5.2.3. DP-based Methods
In DP for FL system, the goal is to ensure the privacy of individual client data during the model training process. In a system with $N$ clients, each client $n$ trains a local model on its own dataset $\mathcal{D}_{n}^{\text{train}}$ by minimizing a local loss function $f_{n}(\cdot)$. After training, the model parameters $\theta_{n}^{t}$ are updated as follows:
(9) $\theta_{n}^{t+1}=\theta_{n}^{t}-\eta\frac{1}{|\mathcal{D}_{n}^{\text{train}}|}\sum_{i=1}^{|\mathcal{D}_{n}^{\text{train}}|}\nabla_{\theta}f_{n}(x_{n,i},y_{n,i}).$
To preserve privacy, each client adds Gaussian noise $\mathcal{N}(0,(\sigma_{n}^{t})^{2}I)$ to the trained local model parameters:
(10) $\tilde{\theta}_{n}^{t+1}=\theta_{n}^{t+1}+\mathcal{N}(0,(\sigma_{n}^{t})^{2}I),$
where $I$ is the identity matrix, and $\sigma_{n}^{t}$ is the noise level at iteration $t$. After adding noise, clients upload their noised model parameters $\tilde{\theta}_{n}^{t+1}$ to the server, which aggregates them using federated averaging:
(11) $\theta_{G}^{t+1}=\frac{1}{N}\sum_{n=1}^{N}\tilde{\theta}_{n}^{t+1}.$
This aggregation ensures that the global model $\theta_{G}^{t+1}$ benefits from the knowledge of all clients while maintaining privacy through the DP mechanism.
5.3. Future Research Directions
Throughout this survey, we have extensively explored the applications of FL in SGs and identified vulnerabilities associated with FL-based SG systems. FL frameworks are commonly utilized across different stages of SGs, including generation, transmission/distribution, and consumption, with the primary goal of preserving privacy. However, there are significant gaps between SOTA FL methodologies and their practical implementation in SGs.
A major challenge is the issue of non-IID data among SG IoT devices. Since patterns vary significantly across devices, traditional FL approaches struggle to maintain model accuracy. Addressing non-IID data requires adaptive techniques that can handle diverse data distributions without compromising performance. Another area of focus is the deployment of large neural networks in FL, which can lead to computational inefficiencies and increased communication overhead. Optimizing the size and efficiency of models in FL contexts is crucial to scaling FL applications in SG environments. Additionally, improving aggregation strategies is essential to ensure effective collaboration between clients. Techniques such as weighted aggregation based on data quality and client reliability can reduce the noise and improve overall model accuracy. Client selection is also an important research direction—intelligent selection mechanisms that prioritize clients with high-quality data and sufficient computational resources can significantly enhance the efficiency and robustness of FL systems in SGs.
On the security front, addressing various types of attacks in FL-based SG systems is paramount. Byzantine attacks, for instance, pose a serious threat as malicious clients can send incorrect model updates, corrupting the global model. These attacks are particularly challenging because they blend in with honest client contributions, making them difficult to detect. Techniques such as robust aggregation methods (e.g., Krum and Bulyan) have been proposed, but further refinement is needed to protect FL systems from Byzantine failures. Backdoor attacks are another prominent threat in which adversaries implant hidden malicious behaviors into the global model while still achieving high performance on regular tasks. This is especially concerning in SG applications, where unnoticed tampering could lead to severe disruptions. Enhancing defense mechanisms such as differential privacy and anomaly detection algorithms can mitigate the risk of backdoor attacks. Inference attacks, where adversaries attempt to extract sensitive information from the shared model, are also a critical area for research. While FL inherently provides some protection by not sharing raw data, adversaries can still exploit the model’s updates to infer private details. Improving privacy-preserving techniques, such as homomorphic encryption and secure multi-party computation, can help protect against these inference attacks. To defend FL-based SG systems from these various threats, several advanced defense mechanisms are being developed. Byzantine-resilient aggregation methods like Trimmed-Mean (trmean_median), Krum (krum), and Bulyan (bulyan) offer robust solutions by filtering out malicious model updates. These approaches, however, need further exploration in the context of SGs, where the scale and complexity of the system introduce unique challenges. FedVRDP (zhang2023byzantine) is another promising approach that combines differential privacy with Byzantine resilience, offering a dual-layered defense against both privacy breaches and malicious attacks. Research into these and other defense strategies will be essential to making FL a viable and secure solution for real-world SG applications.
In conclusion, addressing these challenges and exploring new defense strategies will be critical in advancing FL-based SG systems. The integration of advanced aggregation strategies, efficient model optimization, and robust defense mechanisms will help close the gap between current research and practical applications, ensuring that FL can be securely and efficiently deployed in SG environments.
6. Conclusions
In this paper, we present a comprehensive review of studies focusing on FL-based applications in SGs and their potential vulnerabilities. We begin by comparing our work with existing surveys and outlining our unique contributions. Following this, we define key terminologies related to SGs and FL, providing a brief overview of commonly discussed attack methods within FL frameworks. Next, we examine the deployment of FL in three key stages of SG systems: generation, transmission and distribution, and consumption. The integration of FL at these stages has significantly enhanced data privacy and distribution efficiency in SGs. Furthermore, we analyze and compare the potential vulnerabilities inherent in FL-based SG systems. Finally, we summarize existing research on FL vulnerabilities and propose directions for future research.